Session: May 1999 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A. Cattani at Fri Jun 11 04:49:00 METDST 1999) Success

Effelsberg (A. Kraus at Fri Jun 11 08:51:35 METDST 1999) Minor failures
Problems with the position of tape head 1 occured between UT2235 - 2306 and UT0028 - 0037.

Noto (C.Trigilio at Fri Jun 11 10:42:54 METDST 1999) Minor failures
Due to a problem to the sampler #2, signal from bbc5-6-7-8 sampled at 1-bit. Sampler #2 repaired after the end of the experiment.

Westerbork (Tony Foley at Fri Jun 11 14:43:13 METDST 1999) Success
Used 12 telescopes in array. Playback looked very good.

Torun (K. Borkowski at Fri Jun 11 15:07:07 METDST 1999) Success
No known problems.

Onsala (A. Polatidis at Mon Jun 14 16:42:33 METDST 1999) Success
No known problems

Jodrell (Alastair Gunn at Fri Jun 18 10:57:13 METDST 1999) Success
No known problems.

Calibration (Jean\-Francois Desmurs at Mon Jun 21 16:12:32 METDST 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may99 directory under the name: fr004.antab. Read the comments in the file to get further information.